If the gas has a molar mass of 86 g/mol, what is the density of a gas (in g/L) at 25.0°C and 124.1 kPa?

Consider four identical 1.0-L flasks containing the following gases each at 25°C and 1 atm pressure. For which gas do the molecules have the greatest average kinetic energy?

A.

Ne

B.

Ar

C.

He

D.

Kr

E.

same for all the gases
Answer #1

1)

P= 124.1 kPa

= (124.1*1000) pa

= 124100.0 Pa

P= 124100.0 Pa

= (124100.0/101325) atm

= 1.2248 atm

T= 25.0 oC

= (25.0+273) K

= 298 K

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

1.2248 atm *86.0 g/mol = density * 0.08206 atm.L/mol.K *298.0 K

density = 4.3073 g/L

Answer: 4.31 g/L

2)

Since temperature is same for all the gases, the kinetic energy will also be same.

Answer: E